Find the equation of the line which satisfies each set of conditions:
passes through the point with coordinates (-2, 1) and has slope 2

Answer by venugopalramana(3286)

(-2, 1) and has slope 2
The eqn.of a line with slope m and passing through point x1,y1 is given by
y-y1=m*(x-x1)
y-1=2(x+2)
y-1=2x+4
y=2x+5